# arXiv:2603.00415v1[eess.SY] 28 Feb 2026

### v1

@@ IX-B5 Revisiting SSTs for AI DCs
SSTs [111, 119, 51] are considered primarily for enabling medium-voltage AC to high-voltage DC conversion (e.g., 800 V DC or higher) to improve rack-level power density in AI DCs [52]. In SST-based architectures, traditional UPS functionalities can be embedded within multi-stage power electronic converters. Moreover, SSTs eliminate the distribution transformer in the AI DC power delivery system, thereby minimizing the side effects of bidirectional power flow. However, SSTs face several challenges including lower reliability due to multi-level, multi-stage power conversion, limited power rating (typically a few MVA) and increased control complexity and cost. While SSTs are attractive for high-density power delivery, they are not yet suitable replacements for UPS systems at gigawatt-scale DCs. Multi-level modular SST architectures, as proposed in recent industry studies, may improve scalability, but reliability and standardization remain key challenges.
IX-B6 AI DC as a Microgrid
AI DC microgrids adopt hybrid power architectures [55, 144]. These architectures integrate BESS, conventional generators, distributed generation units, and utility grid connections (see Fig. 8). The objective is to achieve high reliability and redundancy. The placement of BESS is a key design choice. Proper placement allows BESS to damp fast load fluctuations. It also improves local voltage and frequency stability. BESS enable coordinated operation among different power generation units. They support grid services such as demand response and peak shaving. They also provide fast reserves. Optimal integration of BESS with other units requires joint power flow and control design. This integration raises the question of BESS versus grid-interactive UPS systems. The boundary between backup and active grid support becomes unclear. In this context, BESS redefine power interface thresholds and reshape power quality metrics. This enables data center microgrids to act as active and flexible grid participants.
@@ X Conclusions
This paper provides a comprehensive review of multi-layer ESSs in future AI DCs and their role in supporting both DCs and the utility grid. ESSs ranging from chip-level units to large-scale BESSs are discussed, with particular emphasis on their contribution to grid support and ancillary services. Chip-level ESSs act as the first layer for smoothing heterogeneous GPU power profiles, while large-scale BESSs constitute the final layer in this hierarchical structure. The concept of GiUPS is thoroughly examined, highlighting bidirectional power flow capability as a key enabler of grid-aware UPS design. Moreover, the paper extends beyond battery-based ESSs and investigates the roles of FCs and TEs. FCs are considered a promising alternative to conventional gensets in future AI DCs, offering a lower carbon footprint when using hydrogen. The role of TEs in enabling advanced cooling systems and maximizing the utilization of RESs as on-site generation units is also discussed. SLBESSs offer more cost-efficient solutions for ESS integration but require additional monitorization due to lifetime uncertainties. Furthermore, the main challenges associated with ESS integration in AI DCs are outlined. Finally, as a future direction, the deployment of multi-layer ESS architectures necessitates advanced monitoring systems and EMSs to enable optimal coordination across different layers.

### v2

@@ IX-B5 Revisiting Solid-State Transformers (SSTs) for AI DCs
SSTs [109, 117, 50] are primarily considered for enabling medium-voltage AC to high-voltage DC conversion (e.g., 800 V DC or higher) to improve rack-level power density in AI DCs [51]. In SST-based architectures, traditional UPS functionalities can be embedded within multi-stage power electronic converters, and the distribution transformer in the AI DC power delivery system can be eliminated, minimizing side effects of bidirectional power flow [140]. However, SSTs face challenges including lower reliability due to multi-level power conversion stages, limited power ratings (typically a few MVA), and increased control complexity and cost. While attractive for high-density power delivery, SSTs are not yet suitable replacements for UPS systems at gigawatt-scale DCs. Multi-level modular SST architectures may improve scalability, but reliability and standardization remain key challenges requiring further investigation.
IX-B6 Integrated AI DC Microgrid Design and Operation
AI DC microgrids are emerging as a promising architecture for improving both internal power resilience and external grid support [54, 144]. These microgrids adopt hybrid power architectures that combine BESS, conventional generators, distributed generation units, and utility grid connections (see Fig. 9) to achieve high reliability, redundancy, and operational flexibility. Within this framework, the placement and coordination of BESS are critical design choices, since properly located BESS can damp fast AI workload-induced load fluctuations, enhance local voltage and frequency stability, and enable coordinated operation among diverse power generation units.
In addition to supporting internal reliability, BESS can also provide grid services such as demand response, peak shaving, and fast reserves. This broader functionality increasingly blurs the traditional boundary between backup-oriented systems and active grid-support assets, particularly in relation to GiUPS systems, and raises new questions regarding their joint design and control. More broadly, the integration of GiUPS systems, grid-scale BESSs, FC backup generation, TES for cooling, and renewable on-site generation under a unified hierarchical EMS points toward a future in which AI DCs operate as active microgrids capable of seamless transitions between grid-connected and islanded modes. In this vision, AI DCs would no longer behave merely as large and potentially disruptive loads, but rather as flexible grid participants that can reshape power interface thresholds and redefine power quality metrics. However, several open challenges remain, including optimal microgrid topology design for AI DC load characteristics, hierarchical control under uncertainty, and the techno-economic co-optimization of all ESS layers jointly with compute scheduling.
@@ X Conclusions
This paper has provided a comprehensive critical review of multi-layer ESSs for AI DCs and their roles in supporting both facility-level reliability and utility grid integration. Organized around a four-layer hierarchical taxonomy, namely chip-level buffers, rack/server-level ESSs, facility-level GiUPS systems, and grid-scale BESSs, the review systematically analyzes each technology’s response timescale, application role, advantages, limitations, and coordination requirements.
Several key findings emerge from this review. First, AI DC load profiles are fundamentally different from traditional grid loads. Sub-second power variability, high peak-to-average ratios, and unpredictable workload transitions make conventional ESS dispatch strategies insufficient. Second, GiUPS systems represent the most significant architectural evolution in DC power infrastructure, transforming passive UPS batteries into active grid assets capable of FFR, VRT, peak shaving, and synthetic inertia provision. Third, grid-scale BESSs, whether BTM or FTM, are essential for demand charge reduction, renewable integration, and ancillary service provision at the facility level, but their sizing must explicitly account for degradation, load uncertainty, and multi-objective operational requirements. Fourth, non-battery technologies, namely FCs and TES, are underutilized in the current literature relative to their potential. FCs offer a viable clean replacement for diesel backup generation, while TES enables significant cooling efficiency gains and RES utilization. Fifth, SLBESSs offer a cost-effective deployment pathway for large-scale storage in AI DCs, provided that SoH uncertainty and cell heterogeneity are addressed through rigorous degradation-aware control.
The review identifies eight critical open challenges: the absence of multi-physics simulation tools for AI DC EMS validation, the immaturity of energy-aware GPU scheduling, bidirectional power flow impacts on distribution transformers, insufficient load forecasting methods for AI workloads, inadequate degradation models under AI DC cycling conditions, the lack of hierarchical ESS coordination frameworks, LDES technology immaturity, and the absence of multi-layer techno-economic sizing methodologies. Future research should address these challenges through the co-design of compute scheduling and ESS dispatch, the development of high-resolution load monitoring infrastructure, advanced degradation and RUL estimation methods for AI DC operating conditions, and integrated AI DC microgrid architectures. Addressing these open problems will be essential for transforming AI DCs from large disruptive grid loads into flexible, grid-supportive assets that contribute to power system reliability and the sustainable integration of renewable energy.